OpenStudy (anonymous):

factor and solve. 14x² +7x = 0

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf 14x^{2} +7x = 0 }\) factor out of 7x \(\LARGE\color{blue}{ \bf 7x(2x +1) = 0 }\) divide both sides by 7x \(\LARGE\color{blue}{ \bf 2x +1= 0 }\) doit from here.

OpenStudy (anonymous):

hi friend ,. same type quesion again and again ...u beeter try ur self once ... u can check ur ans here .. .. give out ur ans ... we will say wether its rt or wrong ,,,,

OpenStudy (anonymous):

\[\huge 14x² +7x = 0 \rightarrow 7x(2x+1)=0\]

OpenStudy (anonymous):

7x=0 or 2x+1 = 0

OpenStudy (anonymous):

x=0 or x = -1/2

OpenStudy (anonymous):

I'm stuck on the 2x + 1 = 0